OpenStudy (anonymous):

Identify the zeros of the function y=(x+6)(x+6)

OpenStudy (anonymous):

A zero (aka root, solution) of a function is a value (for x in this case) that makes the value of the function ( y in this case) equal to 0

OpenStudy (anonymous):

- Meaning ?

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

(x+6)(x+6)=y (x+6)(x+6)=0 (x+6)=0

OpenStudy (anonymous):

- So does it equal to 6 ?

OpenStudy (anonymous):

What happens if u put 6 into your equation?

OpenStudy (anonymous):

I thought thats what would happen . 6 . ? Idk , im confused .

OpenStudy (anonymous):

x=-6